OpenStudy (chosenmatt):

Simplify the expression. 10 power of 0

OpenStudy (anonymous):

heeeeyyyooo matt!!!

OpenStudy (chosenmatt):

hi

OpenStudy (chosenmatt):

@AriPotta

OpenStudy (anonymous):

Anything raised to the power 0 except 0 is \(1\)..

OpenStudy (anonymous):

\[(x)^0 = 1 . \quad x \ne 0\]

OpenStudy (chosenmatt):

OH YAH I FORGOT

sammixboo (sammixboo):

Haha, everyone forgets something here and there

OpenStudy (anonymous):

And you are happy with that??

OpenStudy (chosenmatt):

well this is my first day studying what im doing...

OpenStudy (anonymous):

wow i was all happy and then you were like (hi) like someone went behind your back and talked like you were never important in their lives :,(

OpenStudy (chosenmatt):

huh?

OpenStudy (anonymous):

Anything wrong with you @RainbowBubbles01 ??

OpenStudy (anonymous):

o im just fine this is the third person in a row who just said hi instead of like hiiiiiii!!!!!! :D

OpenStudy (chosenmatt):

ok then...hiiiiiiiiiiiiiiiiiiiiiiiiii

OpenStudy (anonymous):

hhhhhhhhhhiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiiii @RainbowBubbles01 What's up?

OpenStudy (anonymous):

to late now... :/

OpenStudy (anonymous):

hiiiiiiii water

OpenStudy (anonymous):

Keep smiling friends. :)

OpenStudy (chosenmatt):

Simplify. 2 power of 3 plus 5 power of 2 @waterineyes

OpenStudy (anonymous):

see your the person everyone needs cause your always positive even when something bad happpened :D

OpenStudy (anonymous):

\(2^3 + 5^2\)

OpenStudy (chosenmatt):

yea

OpenStudy (anonymous):

Well, I am always available.. :P

OpenStudy (anonymous):

awesome @waterineyes ever read a tale of two cities?

OpenStudy (anonymous):

\(2^3\) means, multiplying \(2\), three times.. = \(2 \times 2 \times 2 = ??\)

OpenStudy (chosenmatt):

6

OpenStudy (anonymous):

No, busy with Maths all my life. :P But I will..

OpenStudy (chosenmatt):

A. 16 B. 33 C. 35 D. 75

OpenStudy (anonymous):

\(6\) ??

OpenStudy (anonymous):

\(2 \times 2 =??\)

OpenStudy (chosenmatt):

4

OpenStudy (anonymous):

Again multiply 4 with 2.

OpenStudy (chosenmatt):

8

OpenStudy (anonymous):

\[\underline{2 \times 2} \times 2 \implies 4 \times 2 = 8\]

OpenStudy (anonymous):

Then how 6??

OpenStudy (chosenmatt):

oh XD

OpenStudy (anonymous):

\[5^2 = 5 \times 5 = 25\]

OpenStudy (anonymous):

So: \(25 + 8 =??\)

OpenStudy (chosenmatt):

33

OpenStudy (anonymous):

Good.. :)

OpenStudy (chosenmatt):

do you code?

OpenStudy (anonymous):

you mean go by the CoC?

OpenStudy (chosenmatt):

well im coding a website...and im kinda new at it...

OpenStudy (anonymous):

what is it?